OpenStudy (anonymous):

What is the square root of negative 12

OpenStudy (klimenkov):

Use the @Hero 's hint with this: \(\sqrt{ab}=\sqrt a\cdot\sqrt b\).

OpenStudy (anonymous):

=(-12)^1/2 =2(-3)^1/2 =2*3i......where i=(-1)^1/2

OpenStudy (anonymous):

I still don't understand

OpenStudy (anonymous):